OpenStudy (ksaimouli):

simplify x^3-9x/x^2-7x+12

OpenStudy (lalaly):

\[\frac{x(x^2-9)}{(x-3)(x-4)}=\frac{x \cancel{(x-3)}(x+3)}{ \cancel{(x-3)}(x-4)}=\frac{x(x+3)}{(x-4)}\]

OpenStudy (ksaimouli):

thats it

OpenStudy (lalaly):

yeah